A loan shark offers to give you $5,909.00 today, but, in exchange, he wants you to pay him $7,868.00 in one year. What is the rate of interest being charged on the account?

Answer Format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))
Answer #1

Interest rate can be found as under

FV = PV(1+r)^n

7868 = 5909(1+r)^1

=1.3315 = (1+r)

r = 0.3315

ie 33.15%

Alternatively one can also find Interest rate as follows

Interest rate = 7868-5909/5909

=1959/5909

=33.15%